OpenStudy (anonymous):

What is the unit rate per bottle if a pack of 12 bottles of juice sells for $6? $0.50 per bottle $0.72 per bottle $5.00 per bottle $7.20 per bottle

OpenStudy (texaschic101):

12/6 = 1/x (12 bottles to $6 = 1 bottle to x $) cross multiply 12x = 6 x = 6/12 x = 1/2 or 0.50 per bottle check... 12 x 0.50 = 6 6 = 6 (correct)